OpenStudy (anonymous):

Choose the correct simplification of 8x^2(5x + 2x^2 - 3). Answer 40x^4 + 16x^3 - 5x^2 40x^4 - 10x^3 + 5x^2 16x^4 - 40x^3 + 24x^2 16x^4 + 40x^3 - 24x^2

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Distributive property?!?

OpenStudy (anonymous):

not sure.. :l Maybe...

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow a(x + y + z) = ax + ay + a z}\)

OpenStudy (anonymous):

k

Parth (parthkohli):

Umm what do you get?

OpenStudy (anonymous):

I'm lost on this 1 dude :l

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 8x^2(5x + 2x^2 - 3) = 8x^2 \times 5x + 8x^2 \times 2x^2 - 8x^2 \times 3 }\)

OpenStudy (anonymous):

one of the bottum 2

Parth (parthkohli):

I see that it is the last one xD you got those two right

OpenStudy (anonymous):

hey can u give me direct answers for the last ones so i can get working on my english?